# DNTextSpotter: Arbitrary-Shaped Scene Text Spotting via Improved Denoising Training

### v3

@@ 3.2 Query Initialization
$\Delta x_{ij}=\left\{\begin{array}[]{ll}(-1)^{m}\alpha_{ij}D^{x}_{ij}&\text{if}\ \text{positive};\\ (-1)^{m}(\alpha_{ij}+1)D^{x}_{ij}&\text{otherwise}.\end{array}\right.$ (3)
$\Delta y_{ij}=\left\{\begin{array}[]{ll}(-1)^{m}\beta_{ij}D^{y}_{ij}&\text{if}\ \text{positive};\\ (-1)^{m}(\beta_{ij}+1)D^{y}_{ij}&\text{otherwise}.\end{array}\right.$ (4)
where $m$ and $n$ are used to control the direction of the offset, with "positive" indicating the coordinates of the positive part.
@@
Noised Content Queries. We designed Mask Character Sliding (MCS), as shown in Fig. 4, to initialize the noised content queries. For the maximum recognition length $T$, which is equal to the $T$ mentioned above, we perform a sliding operation on the valid characters in the positive part. Specifically, we first determine the number of valid characters $t$, which refers to the number of characters in the input sequence that actually have meaning. Then, we calculate the number of times each valid character should be cloned by performing a division operation $\lfloor\frac{T}{t}\rfloor$, in order to evenly distribute the total length $T$ of the sequence to each valid character. Additionally, since the division of $T$ by $t$ may not be exact, there will be a remainder $k=T-\lfloor\frac{T}{t}\rfloor$ , indicating that there are $k$ additional spaces that need to be allocated. To fairly distribute these extra spaces, we assign one additional clone to each of the first $k$ valid characters in the sequence, ensuring that the allocation for each character is as even as possible. Since this operation visually resembles a character sliding operation, we name it Character Sliding. After sliding the characters, we use a mask operation to control the number of consecutive characters, flipping a portion of the consecutive characters into a background label with a certain probability. After processing the positive part, we add noise to the characters in both the positive part and the negative part, causing these characters to flip to other characters with a probability of $\lambda$. These characters are then transformed into noised content queries after being embedded, with all the characters initialized in the negative part being backgrounds.
@@ Appendix B Details of the Instability Measurement
$W^{i}_{n}=\left\{\begin{array}[]{ll}m,&\text{if }P^{i}_{n}\text{ matches }G_{m}\\ -1,&\text{if }P^{i}_{n}\text{ matches nothing}\end{array}\right.$ (16)

### v4

@@ 3.2 Query Initialization
$\Delta x_{ij}=\left\{\begin{array}[]{ll}(-1)^{u}\alpha_{ij}D^{x}_{ij}&\text{if}\ \text{positive};\\ (-1)^{u}(\alpha_{ij}+1)D^{x}_{ij}&\text{otherwise}.\end{array}\right.$ (3)
$\Delta y_{ij}=\left\{\begin{array}[]{ll}(-1)^{v}\beta_{ij}D^{y}_{ij}&\text{if}\ \text{positive};\\ (-1)^{v}(\beta_{ij}+1)D^{y}_{ij}&\text{otherwise}.\end{array}\right.$ (4)
where $u$ and $v$ are used to control the direction of the offset, with "positive" indicating the coordinates of the positive part.
@@
Noised Content Queries. We designed Mask Character Sliding (MCS), as shown in Fig. 4, to initialize the noised content queries. For the maximum recognition length $T$, which is equal to the $T$ mentioned above, we perform a sliding operation on the valid characters in the positive part. Specifically, we first determine the number of valid characters $t$, which refers to the number of characters in the input sequence that actually have meaning. Then, we calculate the number of times each valid character should be cloned by performing a division operation $\left\lfloor\frac{T}{t}\right\rfloor$, in order to evenly distribute the total length $T$ of the sequence to each valid character. Additionally, since the division of $T$ by $t$ may not be exact, there will be a remainder $k=T-\left\lfloor\frac{T}{t}\right\rfloor$ , indicating that there are $k$ additional spaces that need to be allocated. To fairly distribute these extra spaces, we assign one additional clone to each of the first $k$ valid characters in the sequence, ensuring that the allocation for each character is as even as possible. Since this operation visually resembles a character sliding operation, we name it Character Sliding. After sliding the characters, we use a mask operation to control the number of consecutive characters, flipping a portion of the consecutive characters into a background label with a certain probability. After processing the positive part, we add noise to the characters in both the positive part and the negative part, causing these characters to flip to other characters with a probability of $\lambda$. These characters are then transformed into noised content queries after being embedded, with all the characters initialized in the negative part being backgrounds.
@@ 3.3 Single Attention Mask
$a_{ij}=\left\{\begin{array}[]{ll}1,&\text{ if }j<g\times 2n\text{ and }\left\lfloor\frac{i}{2n}\right\rfloor\neq\left\lfloor\frac{j}{2n}\right\rfloor;\\ 1,&\text{ if }j<g\times 2n\text{ and }i\geq g\times 2n;\\ 0,&\text{otherwise.}\end{array}\right.$ (7)
@@ Appendix B Details of the Instability Measurement
$W^{i}_{n}=\left\{\begin{array}[]{ll}m,&\text{if }P^{i}_{n}\text{ matches }G_{m};\\ -1,&\text{if }P^{i}_{n}\text{ matches nothing.}\end{array}\right.$ (16)